Find a polynomial function of lowest degree with rational coefficients and the following as some of its zeros.
A. –8, 5i
B. 5, − (square root)7, 1 + i

Answer by stanbon(75887)

Find a polynomial function of lowest degree with rational coefficients and the following as some of its zeros.
A. –8, 5i
If 5i is a zero and the coefficients must be rational, -5i is also a zero.
--------------
f(x) = (x+8)(x-5i)(x+5i) = (x+8)(x^2+25)
------------------------------------------------
B. 5, − (square root)7, 1 + i
Same reason as above::
f(x) = (x-5)(x+sqrt(7))(x-sqrt(7))((x-1)-i)((x-1)+i)
----
f(x) = (x-5)(x^2-7)((x-1)^2+1)
----
f(x) = (x-5)(x^2-7)(x^2+2x+2)
--------------------------------
Cheers,
Stan H.
-----------------------